OpenStudy (anonymous):

The half-life of Radium-226 is 1590 years. If a sample contains 300 mg, how many mg will remain after 4000 years?

OpenStudy (anonymous):

\[300\times( \frac{1}{2})^{\frac{4000}{1590}}\]

OpenStudy (anonymous):

another words, the answer is about 37-38 mg

OpenStudy (anonymous):

The way to logically think about sattelite73's response is that the 1/2 is going to be multiplied on 4000/1590 time, which is the number of half-lives in the period that you specified (4000/1590 is amount of time divided by time per half-life). These halves are going to be multiplied onto the original amount 300, so we're halving it the correct number of times. Does that make sense?